# Peer Review Status: Submitted to Seismica

### v1

@@ 1.3 Exact random noise to re-obtain the exponential distribution
Let $M_{i}$ be the geometric random variable with parameter $q=1-e^{\beta\Delta m}$. We now seek to derive the exact distribution for dithering binned magnitudes, such that the exponential distribution with parameter $\beta$ is recovered for the variable $M=M_{i}+Y$.
Since $M_{i}$ creates jumps of size $\Delta m$, the final continuous exponential distribution can be seen as a continuous mixture of geometric steps, dithered by a random offset $Y\in[0,\Delta m)$. It is therefore natural to assume that the dithering variable $Y$ is zero outside this interval, i.e., its support is $0,\Delta m)$. Then, the convolution relating the PDFs of $M,M_{i}$ and $Y$ can be written as
$f_{M}(m)$ $=\sum_{k=1}^{\infty}\mathbb{P}(M_{i}=k)f_{Y}(m-k\Delta m),$ (13)
@@
Remark. Since $Mi$ is defined starting from $i=1$ (see Eq. (1)), while the support of $Y$ is $0,\Delta m)$, it could be more intuitive to select as non-zero term in the sum (13) the right extreme of the interval $k\Delta m,(k+1)\Delta m)$, yielding $f_{M}(m)=\sum_{k=1}^{\infty}\mathbb{P}(M_{i}=k+1)f_{Y}(m-(k+1)\Delta m),$ slightly more complicated calculations and obviously the same result as in (14).
1.3.1 Numerical test
We perform a numerical experiment analogous to that described in Section 1.2.2. In this case, however, the binned magnitudes are dithered by adding a truncated exponential random variable supported on $0,\Delta m)$. The procedure is summarized in Algorithm 2. Results are shown in Table 3 and Figure 4.
Algorithm 2 Lilliefors rejection rate as function of bin width and catalog size - Truncated exponential noise
[…]
3:$b=1$ ( $\beta=b\cdot\ln(10)$ )
[…]
6:for each bin width $\Delta m$ do
7:   for each catalog size $n_{\mathrm{events}}$ do
8:    $counter=0$
9:    for $j=1$ to $N_{\mathrm{SIM}}$ do $\triangleright$ Simulate one catalog
10:     for $i=1$ to $n_{\mathrm{events}}$ do $\triangleright$ Generate binned magnitudes
11:        Draw $k_{i}\in\{0,1,2,\dots\}$ with $\Pr(k_{i}=k)=q(1-q)^{k}$, $q=1-e^{-\beta\Delta m}$
12:        Set $M_{i}=M_{\min}+k_{i}\cdot\Delta m$
13:     end for
14:     Initialize empty list $P$ $\triangleright$ store $p$-values
15:     for $k=1$ to $N_{\mathrm{NOISE}}$ do $\triangleright$ Dither magnitudes
16:        Draw noise
$\varepsilon_{i}=-\frac{1}{\beta}\ln\Bigl(1-U_{i}(1-e^{-\beta\Delta m})\Bigr),\qquad U_{i}\sim\mathcal{U}(0,1),\;i=1,\dots,n_{\mathrm{events}}$
17:        $M_{i}^{\mathrm{dithered}}=M_{i}+\varepsilon_{i}$
18:        Shift magnitudes: $M_{i}^{\text{shifted}}=M_{i}^{\text{dithered}}-M_{\min}$
19:        Apply Lilliefors test: $p_{k}=\mathrm{Lilliefors}(M^{\mathrm{shifted}},\mathrm{dist}=\mathrm{exp})$
20:        Append $p_{k}$ to $P$
21:     end for
22:     Compute mean p-value: $\bar{p}=\mathrm{mean}(P)$
23:     if $\bar{p}<\alpha$ then $counter\leftarrow counter+1$
24:     end if
25:    end for
26:    Rejection rate: $R=100\cdot\frac{counter}{N_{\mathrm{SIM}}}$
27:   end for
28:end for
@@ 1.3.1 Numerical test
Rejection rates tend to increase with catalog size, though at a much slower rate than under the uniform–noise case. This is consistent with the statistical power of Lilliefors test increasing as the catalog size grows, meaning that even very small deviations from the exponential model (e.g., floating rounding, ties, p-value approximation errors) are detected. Rejection rates also increase as $\Delta m$ decreases. This is plausibly related to finite–precision effects: when $\Delta m$ is small, the truncated exponential noise is supported on a very narrow interval, and this could introduce small distortions that the test detects only when the catalog size is very large.
Rejection rates remain below 10% for all combinations of $\Delta m$ and $N$, except for two cases at $N=10^{6}$, where we observe a rejection rate of 11% for $\Delta m=0.2$ and 15% for $\Delta m=0.1$. Nevertheless, these rates are far below those obtained by uniform dithering, and remain close to $\alpha=0.1$ even for $N=10^{6}$. This indicates that dithering magnitudes with truncated exponential noise is more consistent with the assumed GR exponential model.
These conclusions remain unchanged when alternative numerical choices are adopted. In particular, using the median instead of the mean of the $N_{\mathrm{NOISE}}$ p-values (line 22 in Algorithms 1 and 2), or generating binned magnitudes by first sampling from the continuous GR exponential distribution and then discretizing them with the floor function, yields rejection rates that are statistically indistinguishable from those reported here.
@@ 2 Discussion and conclusions
In this study, we show both analytically and numerically that uniform dithering does not correctly recover the underlying continuous exponential distribution. Instead, it introduces a systematic bias that remains undetected at low–to–intermediate catalog sizes due to two opposing effects neutralizing each other: the smoothing introduced by dithering and the limited statistical power of the Lilliefors test. For small and medium catalogs, the uniform noise makes the data ‘appear’ exponential, when they actually are not. As either the bin width increases (= higher approximation error) or the catalog size grows (= higher test’s statistical power), the Lilliefors test becomes sensitive enough to detect this deviation. The result is a systematic overestimation of the magnitude of completeness for large catalogs. As progressively larger candidate values of $M_{c}$ are considered, the number of events with $M\geq M_{c}$ decreases exponentially, thus reducing the effective sample size and, consequently, the statistical power of the Lilliefors test.
To address this issue, we propose replacing uniform dithering with truncated exponential dithering. We demonstrate, both analytically and numerically, that the truncated exponential noise correctly reconstructs a continuous exponential distribution without introducing artifacts into the final magnitude distribution. In principle, one could test the exponentiality hypothesis directly on binned magnitudes using methods like the maximum-likelihood estimation applied to the geometric distribution. Although feasible in principle, this is beyond the scope of this work. Our goal here is to propose a minimal modification to the established practice that completely resolves the issue identified above, while remaining simple to implement.
@@
Code availability
A Python implementation of the the proposed truncated exponential dithering can be found on Zenodo at this link: https://doi.org/10.5281/zenodo.17939386.
Acknowledgements
We thank the Istituto Nazionale di Geofisica e Vulcanologia, Italy, grant "Progetto INGV Pianeta Dinamico” (NEMESIS, NEar real-tiME results of Physical and StatIstical Seismology for earthquakes observations, modeling, and forecasting) - code CUP D53J19000170001 - funded by Italian Ministry MIUR (“Fondo Finalizzato al rilancio degli investimenti delle amministrazioni centrali dello Stato e allo sviluppo del Paese”, legge 145/2018).

### v2

@@ 1.3 Exact random noise to re-obtain the exponential distribution
Let $M_{i}$ be a geometric random variable with parameter $q=1-e^{-\beta\Delta m}$. We now seek to derive the exact distribution for dithering binned magnitudes, such that the exponential distribution with parameter $\beta$ is recovered for the variable $M=M_{i}+Y$.
Since $M_{i}$ creates jumps of size $\Delta m$, the final continuous exponential distribution can be seen as a continuous mixture of geometric steps, dithered by a random offset $Y\in[0,\Delta m)$. It is therefore natural to assume that the dithering variable $Y$ has support confined to this interval, i.e., $Y=0$ outside $0,\Delta m)$. We therefore adopt a left-edge discretization convention, where $M_{i}$ represents the left edge of each magnitude bin of width $\Delta m$. Note that, in Section 1.2, magnitudes were discretized using bin centers, following common practice (see for example the magnitude class representation (1) by tinti:1987). While both conventions are legitimate and mathematically equivalent, the left-edge formulation is more convenient here because it avoids splitting the noise support into two symmetric intervals around the bin center (importantly, the specific discretization convention does not affect the conclusions of this work).
The convolution relating the PDFs of $M,M_{i}$ and $Y$ can be written as
$f_{M}(m)$ $=\sum_{k=1}^{\infty}P(M_{i}=k)f_{Y}(m-k\Delta m),$ (13)
@@
1.3.1 Numerical test: binned magnitudes dithered with truncated exponential noise
We perform two numerical experiments analogous to those described in Section 1.2.2. In the first test, we generate synthetic catalogs of binned magnitudes for increasing catalog sizes, while keeping all other parameters fixed (bin width $\Delta m=0.1$, bb\,-value =1=1, and $M_{\min}=1.0$). This time, however, the binned magnitudes are dithered by adding a truncated exponential random variable supported on $0,\Delta m)$. For each catalog size, the magnitude of completeness $M_{c}$ is calculated as the mean of the Lilliefors test ($\alpha=0.1$) estimates obtained from 50 independent synthetic catalogs. For each synthetic catalog, 100 independent dithering realizations are considered in order to reduce sensitivity to individual realizations of the truncated exponential noise, following herrmann2021inconsistencies. The resulting pp\,-values are averaged and compared to the significance level $\alpha$. This averaged quantity should not be interpreted as a formal pp\,-value combination, but as a way to ensure that the rejection decision does not depend on a specific noise realization.
Table 3 shows that the estimated magnitude of completeness remains stable and close to the true value ($M_{c}\sim M_{\min}=1.0$) across all catalog sizes, including the largest ones.
Catalog Size 1,000 10,000 20,000 50,000 100,000 1,000,000
$M_{c}$ 1.00 1.01 1.01 1.02 1.01 1.03
In the second numerical test, we explore how the rejection probability of the Lilliefors test varies for various combinations of $\Delta m$ and size. The procedure is summarized in Algorithm 2 in the Supplemental Material (Section S5). Results are shown in Table 4 and Figure 4.
@@ 1.3.1 Numerical test: binned magnitudes dithered with truncated exponential noise
Rejection probability tends to increase with catalog size, though at a much slower rate than in the uniform–noise case. This is consistent with the statistical power of Lilliefors test increasing as the catalog size grows, meaning that even very small deviations from the exponential model are detected. These departures can arise from numerical precision effects such as tiny numerical rounding errors in computer calculations, repeated values caused by limited numerical precision, or minor inaccuracies introduced by approximate pp\,-value calculations. Rejection probability also increases as $\Delta m$ decreases. This is plausibly related to finite–precision effects: when $\Delta m$ is small, the support of the truncated exponential noise is very narrow, and this could introduce small distortions that the test detects only when the catalog size is very large.
Rejection probability remains below 10% for all combinations of $\Delta m$ and $N$, except for two cases at $N=10^{6}$, where we observe a rejection probability of 10.6% for $\Delta m=0.2$ and 13.7% for $\Delta m=0.1$. We interpret this behavior as resulting from numerical precision effects (see above) that are present at all sample sizes, but become more evident as the catalog size grows very large. Nevertheless, all rejection probabilities are far below those obtained with uniform dithering, and remain close to $\alpha=0.1$ even for $N=10^{6}$. This indicates that dithering magnitudes with truncated exponential noise is consistent with the assumed GR exponential model.
These conclusions remain unchanged when alternative numerical choices are adopted. In particular, using the median instead of the mean of the $N_{\mathrm{NOISE}}$ pp\,-values (line 22 of Algorithms 1 and 2 in the Supplemental Material), or generating binned magnitudes by first sampling from the continuous GR exponential distribution and then discretizing them with the floor function, yields rejection probabilities that are statistically indistinguishable from those reported here.
@@ 2 Discussion and conclusions
In this study, we show both analytically and numerically that uniform dithering does not correctly recover the underlying continuous exponential distribution. Instead, it introduces a systematic bias that remains undetected at low–to–intermediate catalog sizes due to two opposing effects neutralizing each other: the smoothing introduced by dithering and the limited statistical power of the Lilliefors test. In other words, for small and medium catalogs, the uniform noise makes the data ‘appear’ exponential, when they actually are not. As either the bin width increases (= higher approximation error) or the catalog size grows (= more information), the Lilliefors test becomes sensitive enough to detect this deviation. The result is a systematic overestimation of the magnitude of completeness, with biases exceeding one magnitude unit in large, high-resolution catalogs.
To address this issue, we propose replacing uniform dithering with truncated exponential dithering. We demonstrate, both analytically and numerically, that the truncated exponential noise correctly reconstructs a continuous exponential distribution without introducing artifacts into the final magnitude distribution. One could test the exponentiality hypothesis directly on binned magnitudes using methods like the maximum-likelihood estimation applied to the geometric distribution. Although feasible in principle, this is beyond the scope of this work. Our goal here is to propose a minimal modification to the established practice that completely resolves the issue identified above, while remaining simple to implement.
@@
Acknowledgements.
We thank the Istituto Nazionale di Geofisica e Vulcanologia, Italy, grant "Progetto INGV Pianeta Dinamico” (NEMESIS, NEar real-tiME results of Physical and StatIstical Seismology for earthquakes observations, modeling, and forecasting) - code CUP D53J19000170001 - funded by Italian Ministry MIUR (“Fondo Finalizzato al rilancio degli investimenti delle amministrazioni centrali dello Stato e allo sviluppo del Paese”, legge 145/2018). We thank Davide Zaccagnino and Giuseppe Petrillo for their helpful insights and constructive comments.
Data and code availability
A Python implementation of the proposed truncated exponential dithering can be found on Zenodo at this link: https://doi.org/10.5281/zenodo.17939386.
Competing interests
The authors acknowledge that there are no conflicts of interest recorded.